Dr. Woz and Dr. Yaz want to test whether violent video games have an effect on aggressive thinking through the use of a One Sample t test. In order to test aggressive thinking, Woz and Yaz are going to use a test developed at GTA University; scores greater than 70 on the test indicate agreement with statements that endorse aggressive behavior. They select a sample of n = 100 children from a local primary school for the study. Each of the students spends an hour playing a first-person shooter-type video game and then takes the test. The mean and standard deviation for the sample is M = 73 and s = 20. Woz and Yaz would like to know if the sample mean of 73 is statistically significantly different from 70 using a two-tailed test with α < .05 and critical values of t of ±1.98.  

What is the Null Hypothesis?

What is the Research Hypothesis?

What is the estimated standard error?

Calculate the t value for the sample mean

Should the null be accepted or rejected?

Explain your reasoning for your decision regarding the Null Hypothesis

Write up your results in the proper APA format
Answer #1

Ans:

degrees of freedom=100-1=99

As,critical t value=+/-1.98,rejection region is: t<-1.98 or t>1.98

Null Hypothesis:

Research Hypothesis

estimated standard error=20/sqrt(100)=2

Test statistic:

t=(73-70)/2

t=1.5

As,test statistic does not fall in rejection region,we do not reject the null hypothesis(i..e we accept the null hypothesis)

The sample mean of 73 is not significantly different from 70.

t(99)=1.50,p>0.05,not significant